# LeanGaussian: Breaking Pixel or Point Cloud Correspondence in Modeling 3D Gaussians

### v2

@@ Abstract
Rencently, Gaussian splatting has demonstrated significant success in novel view synthesis. Current methods often regress Gaussians with pixel or point cloud correspondence, linking each Gaussian with a pixel or a 3D point. This leads to the redundancy of Gaussians being used to overfit the correspondence rather than the objects represented by the 3D Gaussians themselves, consequently wasting resources and lacking accurate geometries or textures. In this paper, we introduce LeanGaussian, a novel approach that treats each query in deformable Transformer as one 3D Gaussian ellipsoid, breaking the pixel or point cloud correspondence constraints. We leverage deformable decoder to iteratively refine the Gaussians layer-by-layer with the image features as keys and values. Notably, the center of each 3D Gaussian is defined as 3D reference points, which are then projected onto the image for deformable attention in 2D space. On both the ShapeNet SRN dataset (category level) and the Google Scanned Objects dataset (open-category level, trained with the Objaverse dataset), our approach, outperforms prior methods by approximately 6.1%, achieving a PSNR of 25.44 and 22.36, respectively. Additionally, our method achieves a 3D reconstruction speed of 7.2 FPS and rendering speed 500 FPS. The code will be released at https://github.com/jwubz123/DIG3D.
@@ 1 Introduction
Mainstream 3D Gaussian modeling methods falls into two primary categories as shown in fig. 1: the first involves direct feature lifting from individual pixels to 3D space [44, 40, 4, 45, 50, 47], where each pixel is associated with one Gaussian ellipsoid; the second method utilizes point clouds and triplane features [58, 49], treating each Gaussian as a point alongside its corresponding triplane feature. However, both representations require alignment with other forms, be it 2D pixels or 3D point clouds. This requirement leads to redundant Gaussians being utilized to essentially overfit other representations rather than focusing on the 3D Gaussians themselves. This clashes with the Gaussian splatting process since the rendered pixels by Gaussian splatting has no one-to-one correspondence to the 3D Gaussians. For instance, as demonstrated in fig. 1, Splatter Image [44], the foremost pixel-aligned method, heavily concentrating (over half) of 3D Gaussians in a single plane (the background). Triplane-Gaussian [58], the most influenced method combining point clouds with triplane features, complicates the modeling process by regressing dense point clouds (fig. 1 (a)). Furthermore, the triplane representation compresses 3D space, leading to a lack of detailed information in the 3D structure and imposing a rigid grid alignment that limits flexibility [45, 36]. Additionally, as illustrated in fig. 1 (b), the computational complexity of Splatter Image increases quadratically with the input resolution due to pixel alignment regression, while Triplane-Gaussian has significantly higher memory costs because of its dense point cloud and complex Transformer design across triplane.
To address the overfitting resulting from pixel or point cloud correspondence, we present LeanGaussian, a novel approach that breaks free from the correspondence constraint by directly modeling 3D Gaussians. Inspired by the effective and efficient Deformable-DETR (Detection Transformer) framework in detection tasks [57, 3, 53, 18, 30, 20, 19], where ‘queries’ are utilized to iteratively refine the regression of bounding box (bbox) parameters (position, size and category), we leverage queries to regress and refine the 3D Gaussian parameters following the DETR paradigm, using image features as keys and values. In our approach, each query represents a Gaussian ellipsoid (or its refinement) with a linear transformation. Queries representing all 3D Gaussians are derived through a layer-by-layer refinement manner within the Transformer decoder. The flexibility to adjust the number of queries arbitrarily ensures that the modeling approach avoids overfitting to alternative representations, whether in pixel or point cloud form.
@@ 3 Methods
In this section, we introduce our proposed LeanGaussian for novel view synthesis (NVS) from single-view images. We begin by providing the necessary preliminaries for 3D Gaussian splatting, as detailed in section 3.1. Next, we give the formulation of the novel view synthesis (NVS) tasks using 3D Gaussian splatting in section 3.2. In section 3.3, we describe the overall framework of our method and give the detailed design of updating 3D Gaussians with deformable Transformer. Finally, we discuss the training objective in section 3.4. The high level difference between our model and Splatter Image is in the section A.1.
@@ 3.3 LeanGaussian
As depicted in fig. 2 (c), within each layer, the queries $\mathbf{q}$ are first updated by a deformable cross attention $\mathrm{GaussianDFA}$ with the projected reference points in section 3.3. To be more precise, the attention scores in the $l$-th layer for the sampling points $\mathbf{S}^{l}$ are computed using Multilayer Perceptron (MLP) on queries, while the values are obtained by bilinearly projecting the image feature at the sampling points. The update of queries cooresponding to the $n$-th Gaussian at the $l$-th layer from $\mathbf{q}_{n}^{l}$ to $\mathbf{q}_{n}^{l^{\prime}}$ through deformable cross-attention is shown in Equation eq. 4.
@@
Subsequently, a self-attention operation $\mathrm{SelfAttn}$ is employed to enable information exchange among all ellipsoids, followed by a feed-forward network $\mathrm{FFN}$. With $\mathbf{P}^{l}$ denoting the reference points at layer $l$ with detail in section 3.3, the decoder layer can be described as eq. 6:
@@
Taking inspiration from [53], we employ multi-layer refinement to enhance the decoding process and facilitate iterative improvement. Firstly, we obtain the initialized 3D Gaussians $\mathbf{G}_{\mathrm{init}}$ in section 3.3 and proceed to train a parameter increment $\mathbf{\Delta G}$ in section 3.3 for each layer. We then update the 3D Gaussian parameters in a sequential manner, rendering and supervising them at each layer. This refinement process is denoted as $\mathbf{\Delta G}^{l}=\mathcal{S}^{l}(\mathbf{q}^{l})$, where $\mathcal{S}^{l}$ represents the splat head of the $l$-th layer. To update the 3D Gaussian parameters from Gaussians $\mathbf{G}$ in the previous layer, we apply the following change: $\mathbf{G}^{l+1}=\mathbf{G}^{l}\bigoplus\mathbf{\Delta G}^{l}$. Here, $\bigoplus$ denotes the update operation of $\mathbf{G}$ by $\mathbf{\Delta G}$, involving multiplication for rotation and addition for other parameters.
@@ 4 Experiments
We assess the effectiveness of our method in NVS using single-view RGB image in this section. The detailed experiment settings are in section A.2. The dataset and evaluation metrics employed in our study, followed by an explanation of our experimental setup. The qualitative and quantitative comparison of novel view synthesis between our method and the previous approaches are in section 4.1. Tthe nessaesary ablation studies for our model design are in section 4.2 More visualization and experimental details can be found in section A.2.
@@
In our study, we employ two datasets, one at the category level and the other at the open-category level. The category-level dataset, ShapeNet-SRN [42], consists of two object classes: ’Cars’ and ’Chairs’. The open-category dataset, Objaverse LVIS [7], which filters the high quality data from the Objaverse with 1156 object categories commonly encountered in daily life. We evaluate the model trained on Objaverse using the Google Scanned Objects (GSO) benchmark [8], which includes 1,030 3D objects across 17 categories. For GSO, we utilize rendered images from Free3D [56], following the dataset splitting methods outlined in zero-1-to-3 [28] and Free3D [56]. Further details about the datasets are in section A.2.
@@ 4.1 Comparison with SOTA methods
For each object, we compute the average metrics (PSNR, SSIM, and LPIPS) for the rendered images in the dataset except for the input image. Our method surpasses all the methods presented in the table. table 1 showcase the results of training on the Objaverse LVIS dataset [7] and test on the GSO dataset [8], our model outperforms previous methods. Particularly, our model achieves a higher PSNR score of 22.36, surpassing previous methods by 1.28. Moreover, Triplane-Gaussian requires 3D supervision while other methods do not. On another note, for the testing of Triplane-Gaussian, we use the lightweight checkpoint provided training under Objaverse-LVIS dataset. In the results outlined in table 2 on the ShapeNet SRN dataset, our model also achieves State-of-the-Art (SoTA) results in both the chairs and cars categories.
@@
To showcase the visual quality of our approach, we present visualizations of the generated novel views on both GSO and ShapNet SRN in fig. 4 and fig. 5, respectively. In the Splatter Image, an additional pizza is observed at the back of the plate, and the bowl exhibits an incorrect coloration. Splatter Image frequently predicts incorrect geometries, whereas Triplane-Gaussian often exhibits inaccuracies in texture rendition. Furthermore, we provide a comparison with generation-based models such as LGM [45] and InstantMesh [49] in section A.3 fig. 9. For further insights, additional visualizations, including a quality comparison between Gamba [41] and our model, can be found in section A.3. We also test the model for the in-the-wild figures and provide the visualization in fig. 11. Additionally, we include point cloud representations (center of Gaussians) to compare the 3D geometry of the objects in section A.3.
@@
Moreover, We specifically choose views with an elevation or azimuth difference exceeding 45 degrees compared to the input view across all objects in the GSO dataset. Subsequently, we compute the mean results across all views for all objects in the GSO dataset. The result is shown in table 3. Our model’s performance remains similar (PSNR reduced by 0.35dB), whereas previous methods experience a significant decline (PSNR reduced by around 1.5dB). We also provide the result in ShapeNet SRN Chair dataset in section A.3.
@@
We conducted speed tests for inference on a single A100 GPU. As depicted in table 4, our inference process is rapid (less than 1s) due to the fast rendering speed (0.002s) of Gaussian splatting. Additionally, leveraging deformable attention, our 3D reconstruction speed (0.14s) surpasses that of other heavy Transformer-based models like the one in [13] (2.24s) or in [58] (1.28s). Additionally, even Triplane Gaussian here is the lightweight version, its dense Transformer and complex triplane design makes the inference slow. For the methods like LGM [45] that first use generation models to generate multiple views from the single input view, the inference speed is much more slow. During inference, our model executes a single forward pass to obtain the 3D Gaussian parameters and multiple passes to generate novel views. Our rendering speed is faster compared to Splatter Image because fewer Gaussians are needed to represent 3D objects. Furthermore, by utilizing a fixed number of Gaussians and disregarding image resolution, our method maintains an almost constant rendering time and slow memory cost change as resolution increases, in contrast to previous pixel-aligned methods that suffer from quadratic growth. This comparison is detailed in fig. 12.
@@ 4.2 Ablation study
We offer a comprehensive ablation study detailing the specifics of our design in table 5. As depicted in the table, supervision in the initial stage enhances the PSNR by approximately 1.5. Utilizing the 3D Gaussians from the first stage as the initialization for the second stage, rather than Gaussians from random queries, yields minimal differences in the results. However, without the deformable decoder, regressing 3D Gaussians from models like DINOv2 or UNet (which has a per-pixel alignment or per-patch (several pixels) alignment) may underperform compared to our model, as shown in table 7.
@@
In table 5, substituting bounding boxes with Gaussians in the deformable Transformer’s decoder yields poor results. Removing multi-layer refinement prevents positive contributions from each layer. Without the 3D reference point, random image feature points may lack accuracy compared to 3D Gaussian positions. Omitting the UNet feature results in sole reliance on the pre-trained model, potentially unsuitable for our task. Without the deformable Transformer, regressing 3D Gaussians from models like DINOv2 or UNet may underperform compared to our model, as shown in section A.3. Furthermore, we also do ablation study on the hyperparameter selections and other model designs with the detailed results in section A.3.
@@ A.3 More results
To prove that our method is not a shortcut for the input view, we conducted an experiment where we removed the views that were very close to the input view. Specifically, we excluded the nearest 8 views (index 60 to 68), as compared to Splatter Image. The results of this experiment, presented in Table table 6, clearly indicate the superiority of our method over Splatter Image.
Our method exhibits a substantial improvement in performance compared to Splatter Image .When comparing table 2 and table 6, our model shows minimal decrease in metrics when removing the values of the 8 views near the input view. However, the Splatter Image dataset exhibits a notable decrement in performance. This comparison provides evidence that our method is not simply a shortcut around the input view. By removing the values of the 8 views near the input view, our method decrease not much, which highlight the effectiveness of our method in enhancing the quality of the generated views when compared to Splatter Image. These results validate that our approach performs better, particularly when it comes to generating novel views that far from input views.
@@
To analyze the impact of the number of Gaussians used for object representation, we conducted an ablation study. Intuitively, a higher number of Gaussians tends to enhance performance but at the cost of slower speed. Thus, we aimed to find a trade-off between model complexity and performance. The results of this study are presented in table 8, indicating that increasing the number of Gaussians leads to improved performance. However, it is crucial to note that using an excessively large number of Gaussians significantly increases the training cost. To provide a more explicit comparison between inference time and PSNR for different numbers of Gaussians, we present fig. 12 (a). When the number of Gaussians smaller 10,000, the PSNR increases significantly. However, when the number of Gaussians exceeds 10,000, the additional benefits diminish while incurring substantial time consumption. Consequently, to strike a balance between speed and quality, we ultimately opted for 10,000 Gaussians to represent an object. This choice allows us to achieve satisfactory results while maintaining reasonable training efficiency.
@@
The comparison between Triplane-Gaussian, Tripo [47] and our methods is shown in table 9. Triplane Gaussian requires 3D supervision and takes longer inference time while get worse performance comparing to our model. We test on the given light-weight checkpoint in the github. Our model surpass the previous methods on both the performance and the inference speed.

### v4

@@ Abstract
Recently, Gaussian splatting has demonstrated significant success in novel view synthesis. Current methods often regress Gaussians with pixel or point cloud correspondence, linking each Gaussian with a pixel or a 3D point. This leads to the redundancy of Gaussians being used to overfit the correspondence rather than the objects represented by the 3D Gaussians themselves, consequently wasting resources and lacking accurate geometries or textures. In this paper, we introduce LeanGaussian, a novel approach that treats each query in deformable Transformer as one 3D Gaussian ellipsoid, breaking the pixel or point cloud correspondence constraints. We leverage deformable decoder to iteratively refine the Gaussians layer-by-layer with the image features as keys and values. Notably, the center of each 3D Gaussian is defined as 3D reference points, which are then projected onto the image for deformable attention in 2D space. On both the ShapeNet SRN dataset (category level) and the Google Scanned Objects dataset (open-category level, trained with the Objaverse dataset), our approach, outperforms prior methods by approximately 6.1%, achieving a PSNR of 25.44 and 22.36, respectively. Additionally, our method achieves a 3D reconstruction speed of 7.2 FPS and rendering speed 500 FPS. Codes are available at https://github.com/jwubz123/LeanGaussian.
@@ 1 Introduction
Mainstream 3D Gaussian modeling methods falls into two primary categories as shown in Fig. 1: the first involves direct feature lifting from individual pixels to 3D space [44, 40, 4, 45, 50, 47], where each pixel is associated with one Gaussian ellipsoid; the second method utilizes point clouds and triplane features [58, 49], treating each Gaussian as a point alongside its corresponding triplane feature. However, both representations require alignment with other forms, be it 2D pixels or 3D point clouds. This requirement leads to redundant Gaussians being utilized to essentially overfit other representations rather than focusing on the 3D Gaussians themselves. This clashes with the Gaussian splatting process since the rendered pixels by Gaussian splatting has no one-to-one correspondence to the 3D Gaussians. As demonstrated in Fig. 1, Splatter Image [44], the foremost pixel-aligned method, heavily concentrates (over half) 3D Gaussians in a single plane (the background). Triplane-Gaussian [58], the most influenced method combining point clouds with triplane features, complicates the modeling process by regressing dense point clouds (Fig. 1 (a)). Furthermore, the triplane representation compresses 3D space, leading to a lack of detailed information in the 3D structure and imposing a rigid grid alignment that limits flexibility [45, 36]. Additionally, as illustrated in Fig. 1 (b), the computational complexity of Splatter Image increases quadratically with the input resolution due to pixel alignment regression, while Triplane-Gaussian has significantly higher memory costs because of its dense point cloud and complex Transformer design across triplane.
To address the overfitting resulting from pixel or point cloud correspondence, we present LeanGaussian, a novel approach that breaks free from the correspondence constraint by directly modeling 3D Gaussians. Inspired by the effective and efficient Deformable-DETR (Detection Transformer) framework in detection tasks [57, 3, 53, 18, 31, 20, 19], where ‘queries’ are utilized to iteratively refine the regression of bounding box (bbox) parameters (position, size and category), we leverage queries to regress and refine the 3D Gaussian parameters following the DETR paradigm, using image features as keys and values. In our approach, each query represents a Gaussian ellipsoid with a linear transformation. Queries representing all 3D Gaussians are derived through a layer-by-layer refinement manner within the Transformer decoder. The flexibility to adjust the number of queries arbitrarily ensures that the modeling approach avoids overfitting to alternative representations, whether in pixel or point cloud form.
@@ 3 Methods
In this section, we introduce our proposed LeanGaussian for novel view synthesis (NVS) from single-view images. We begin by providing the necessary preliminaries for 3D Gaussian splatting, as detailed in Sec. 3.1. Next, we give the formulation of the novel view synthesis (NVS) tasks using 3D Gaussian splatting in Sec. 3.2. In Sec. 3.3, we describe the overall framework of our method and give the detailed design of updating 3D Gaussians with deformable Transformer. Finally, we discuss the training objective in Sec. 3.4. The high level difference between our model and Splatter Image is in the Sec. A.1.
@@ 3.3 LeanGaussian
As depicted in Fig. 2 (c), within each layer, the queries $\mathbf{q}$ are first updated by a deformable cross attention $\mathrm{GaussianDFA}$ with the projected reference points in Sec. 3.3. To be more precise, the attention scores in the $l$-th layer for the sampling points $\mathbf{S}^{l}$ are computed using Multilayer Perceptron (MLP) on queries, while the values are obtained by bilinearly projecting the image feature at the sampling points. The update of queries cooresponding to the $n$-th Gaussian at the $l$-th layer from $\mathbf{q}_{n}^{l}$ to $\mathbf{q}_{n}^{l^{\prime}}$ through deformable cross-attention is shown in Equation Eq. 4.
@@
Subsequently, a self-attention operation $\mathrm{SelfAttn}$ is employed to enable information exchange among all ellipsoids, followed by a feed-forward network $\mathrm{FFN}$. With $\mathbf{P}^{l}$ denoting the reference points at layer $l$ with detail in Sec. 3.3, the decoder layer can be described as Eq. 6:
@@
Taking inspiration from [53], we employ multi-layer refinement to enhance the decoding process and facilitate iterative improvement. Firstly, we obtain the initialized 3D Gaussians $\mathbf{G}_{\mathrm{init}}$ in Sec. 3.3 and proceed to train a parameter increment $\mathbf{\Delta G}$ in Sec. 3.3 for each layer. We then update the 3D Gaussian parameters in a sequential manner, rendering and supervising them at each layer. This refinement process is denoted as $\mathbf{\Delta G}^{l}=\mathcal{S}^{l}(\mathbf{q}^{l})$, where $\mathcal{S}^{l}$ represents the splat head of the $l$-th layer. To update the 3D Gaussian parameters from Gaussians $\mathbf{G}$ in the previous layer, we apply the following change: $\mathbf{G}^{l+1}=\mathbf{G}^{l}\bigoplus\mathbf{\Delta G}^{l}$. Here, $\bigoplus$ denotes the update operation of $\mathbf{G}$ by $\mathbf{\Delta G}$, involving multiplication for rotation and addition for other parameters.
@@ 4 Experiments
We assess the effectiveness of our method in NVS using single-view RGB image in this section. The detailed experiment settings are in Sec. A.2. The dataset and evaluation metrics employed in our study, followed by an explanation of our experimental setup. The qualitative and quantitative comparison of novel view synthesis between our method and the previous approaches are in Sec. 4.1. Tthe nessaesary ablation studies for our model design are in Sec. 4.2 More visualization and experimental details can be found in Sec. A.2.
@@
In our study, we employ two datasets, one at the category level and the other at the open-category level. The category-level dataset, ShapeNet-SRN [42], consists of two object classes: ’Cars’ and ’Chairs’. The open-category dataset, Objaverse LVIS [7], which filters the high quality data from the Objaverse with 1156 object categories commonly encountered in daily life. We evaluate the model trained on Objaverse using the Google Scanned Objects (GSO) benchmark [8], which includes 1,030 3D objects across 17 categories. For GSO, we utilize rendered images from Free3D [56], following the dataset splitting methods outlined in zero-1-to-3 [28] and Free3D [56]. Further details about the datasets are in Sec. A.2.
@@ 4.1 Comparison with SOTA methods
For each object, we compute the average metrics (PSNR, SSIM, and LPIPS) for the rendered images in the dataset except for the input image. Our method surpasses all the methods presented in the table. Tab. 1 showcase the results of training on the Objaverse LVIS dataset [7] and test on the GSO dataset [8], our model outperforms previous methods. Particularly, our model achieves a higher PSNR score of 22.36, surpassing previous methods by 1.28. Moreover, Triplane-Gaussian requires 3D supervision while other methods do not. On another note, for the testing of Triplane-Gaussian, we use the lightweight checkpoint provided training under Objaverse-LVIS dataset. In the results outlined in Tab. 2 on the ShapeNet SRN dataset, our model also achieves State-of-the-Art (SoTA) results in both the chairs and cars categories.
@@
To showcase the visual quality of our approach, we present visualizations of the generated novel views on both GSO and ShapNet SRN in Fig. 4 and Fig. 5, respectively. In the Splatter Image, an additional pizza is observed at the back of the plate, and the bowl exhibits an incorrect coloration. Splatter Image frequently predicts incorrect geometries, whereas Triplane-Gaussian often exhibits inaccuracies in texture rendition. Furthermore, we provide a comparison with generation-based models such as LGM [45] and InstantMesh [49] in Sec. A.3 Fig. 9. For further insights, additional visualizations, including a quality comparison between Gamba [41] and our model, can be found in Sec. A.3. We also test the model for the in-the-wild figures and provide the visualization in Fig. 11. Additionally, we include point cloud representations (center of Gaussians) to compare the 3D geometry of the objects in Sec. A.3.
@@
Moreover, We specifically choose views with an elevation or azimuth difference exceeding 45 degrees compared to the input view across all objects in the GSO dataset. Subsequently, we compute the mean results across all views for all objects in the GSO dataset. The result is shown in Tab. 3. Our model’s performance remains similar (PSNR reduced by 0.35dB), whereas previous methods experience a significant decline (PSNR reduced by around 1.5dB). We also provide the result in ShapeNet SRN Chair dataset in Sec. A.3.
@@
We conducted speed tests for inference on a single A100 GPU. As depicted in Tab. 4, our inference process is rapid (less than 1s) due to the fast rendering speed (0.002s) of Gaussian splatting. Additionally, leveraging deformable attention, our 3D reconstruction speed (0.14s) surpasses that of other heavy Transformer-based models like the one in [13] (2.24s) or in [58] (1.28s). Additionally, even Triplane Gaussian here is the lightweight version, its dense Transformer and complex triplane design makes the inference slow. For the methods like LGM [45] that first use generation models to generate multiple views from the single input view, the inference speed is much more slow. During inference, our model executes a single forward pass to obtain the 3D Gaussian parameters and multiple passes to generate novel views. Our rendering speed is faster compared to Splatter Image because fewer Gaussians are needed to represent 3D objects. Furthermore, by utilizing a fixed number of Gaussians and disregarding image resolution, our method maintains an almost constant rendering time and slow memory cost change as resolution increases, in contrast to previous pixel-aligned methods that suffer from quadratic growth. This comparison is detailed in Fig. 12.
@@ 4.2 Ablation study
We offer a comprehensive ablation study detailing the specifics of our design in Tab. 5. As depicted in the table, supervision in the initial stage enhances the PSNR by approximately 1.5. Utilizing the 3D Gaussians from the first stage as the initialization for the second stage, rather than Gaussians from random queries, yields minimal differences in the results. However, without the deformable decoder, regressing 3D Gaussians from models like DINOv2 or UNet (which has a per-pixel alignment or per-patch (several pixels) alignment) may underperform compared to our model, as shown in Tab. 7.
@@
In Tab. 5, substituting bounding boxes with Gaussians in the deformable Transformer’s decoder yields poor results. Removing multi-layer refinement prevents positive contributions from each layer. Without the 3D reference point, random image feature points may lack accuracy compared to 3D Gaussian positions. Omitting the UNet feature results in sole reliance on the pre-trained model, potentially unsuitable for our task. Without the deformable Transformer, regressing 3D Gaussians from models like DINOv2 or UNet may underperform compared to our model, as shown in Sec. A.3. Furthermore, we also do ablation study on the hyperparameter selections and other model designs with the detailed results in Sec. A.3.
@@ A.3 More results
To prove that our method is not a shortcut for the input view, we conducted an experiment where we removed the views that were very close to the input view. Specifically, we excluded the nearest 8 views (index 60 to 68), as compared to Splatter Image. The results of this experiment, presented in Table Tab. 6, clearly indicate the superiority of our method over Splatter Image.
Our method exhibits a substantial improvement in performance compared to Splatter Image .When comparing Tab. 2 and Tab. 6, our model shows minimal decrease in metrics when removing the values of the 8 views near the input view. However, the Splatter Image dataset exhibits a notable decrement in performance. This comparison provides evidence that our method is not simply a shortcut around the input view. By removing the values of the 8 views near the input view, our method decrease not much, which highlight the effectiveness of our method in enhancing the quality of the generated views when compared to Splatter Image. These results validate that our approach performs better, particularly when it comes to generating novel views that far from input views.
@@
To analyze the impact of the number of Gaussians used for object representation, we conducted an ablation study. Intuitively, a higher number of Gaussians tends to enhance performance but at the cost of slower speed. Thus, we aimed to find a trade-off between model complexity and performance. The results of this study are presented in Tab. 8, indicating that increasing the number of Gaussians leads to improved performance. However, it is crucial to note that using an excessively large number of Gaussians significantly increases the training cost. To provide a more explicit comparison between inference time and PSNR for different numbers of Gaussians, we present Fig. 12 (a). When the number of Gaussians smaller 10,000, the PSNR increases significantly. However, when the number of Gaussians exceeds 10,000, the additional benefits diminish while incurring substantial time consumption. Consequently, to strike a balance between speed and quality, we ultimately opted for 10,000 Gaussians to represent an object. This choice allows us to achieve satisfactory results while maintaining reasonable training efficiency.
@@
The comparison between Triplane-Gaussian, Tripo [47] and our methods is shown in Tab. 9. Triplane Gaussian requires 3D supervision and takes longer inference time while get worse performance comparing to our model. We test on the given light-weight checkpoint in the github. Our model surpass the previous methods on both the performance and the inference speed.